Nobeyama Radiospectrographs (NoRS) Analysis Manual ver. 0.0

T. Yokoyama

July 7, 2000

1 Introduction

This is a manual for analysis of the data obtained by Nobeyama Radiospectrographs (NoRS; stop operation in Aug 1994; Kai et al. 1980).

For any questions and requests, please send an e-mail to

     nsro-service@solar.nro.nao.ac.jp

The latest information on NoRS is updated on NSRO Web site. The URL is

      http://solar.nro.nao.ac.jp/nors/

2 How to setup

(1) Installation of the SolarSoftware.

Install the SolarSoftware (SSW). If this was not installed, please contact your system manager. The primary distribution site for SolarSoftware is: http://www.lmsal.com/solarsoft/

(2) Setup of your personal environment Include the followings in your environment setup file (~/.cshrc). setenv SSW SSW-

directory^{1}

      setenv SSW INSTR ``nors''
      setenv NORS ${SSW}/radio/nors
      source ${SSW}/gen/setup/setup.ssw
      source ${NORS}/setup/setup.nors

Note that the enviroment variable ${SSW} can be different (Ask your system manager)^{2}. In case you analyze many kinds of data at the same time, define SSW INSTR as follows, e.g.

      setenv SSW INSTR ``nors sxt norh''

3 Where is the Data ?

3.1 Raw Data

Some raw NoRS data are put in the NSRO archive. You can obtain them by anonymous FTP.

1/sgi1/ssw in NSRO

2In NSRO, it is /sgi1/ssw

4 Analysis

4.1 Start Analysis

The NoRS IDL procedures described in this section all depend on the SolarSoftware (SSW). When you start the analysis, start up the SSW/IDL as follows:

unix% sswidl <CR>

5 Read the Data

In order to read the data into IDL session, give the file name as follows:

IDL> file=’./ds940220’

IDL> nors rd dat,file,time,mvd,freq,datalo,datahi,mvdd2,freqd2,datad2 <CR> In order to read the data in specified duration:

IDL> timerange=[’1994-2-20 2:00’,’1994-2-20 4:00’] <CR>

IDL> nors rd dat,file,time,mvd,freq,datalo,datahi ,mvdd2,freqd2,datad2,timerange=timerange <CR>

6 Plot

For plotting,

IDL> nors plot,time,mvd,freq,bytscl(dataloi) <CR> For plotting in specified duration

IDL> nors plot,time,mvd,freq,bytscl(dataloi),timerange=timerange <CR>

A NoRS data archive: anonymous FTP

The Nobeyama Solar Radio Observatory (NSRO/NAOJ) has prepared the anonymous FTP server for the NoRS data archive. The URL is

   ftp://nsro-archive.nro.nao.ac.jp/pub/nors

If you need to transfer a large amount of data , please contact

   nsro-service@solar.nro.nao.ac.jp

We will make and mail a CD-ROM for you. The network capability in Nobeyama is limited.

B Acknowledgement

We thank the following people. SolarSoftware is a software package for analysis of Solar Data on IDL. That is supported under various NASA contracts (SXT/EIT/MDI/TRACE/YPOP etc.) and is all in the public domain. IDL is a product by Research System Inc.

References

Kai et al., ”Nobeyama Radiospectrographs for Solar Observations”, 1980, PASJ, 32, 371